OpenStudy (anonymous):

Factor by grouping : 3bc-6bd+2c-4d

myininaya (myininaya):

(3bc-6bd)+(2c-4d) = 3b(c-2d)+2(c-2d) = (c-2d)(3b+2)

OpenStudy (anonymous):

Thank you :)

myininaya (myininaya):

np :)